OpenStudy (anonymous):

What is the current real value of $1000 after 7 years if the annual rate of inflation is 6%? Round to nearest dollar.

OpenStudy (anonymous):

(I'm not sure about this one, but I'll give it a shot!) If each year, the value of goods is increasing by 6%, then each year, the dollar is 0.94 times what it was last year. New value of dollar = previous value of dollar * (0.94)^(number of years) V = v*(0.94)^t V = 1000*(0.94)^7 V = 648.48 So the current real value is $648.

OpenStudy (anonymous):

my answer choices are 665 764 742 150 67

OpenStudy (anonymous):

i would said 665 cause it was close :D

OpenStudy (anonymous):

Me too. (: